OpenStudy (anonymous):

s=r*theta 4in. =7in.*theta Find theta?

OpenStudy (mathmale):

Ms. Garcia: simply divide both sides of this equation by 7 in. After reducing the result, you'll have the angle, theta, in radians. That's it.

OpenStudy (anonymous):

so i divide 4 by 7 & 7 by 7?

OpenStudy (anonymous):

\[4in=7in*\theta\]Divide both sides by 7in.\[\frac{4in}{7in}=\theta\]Inches divided by inches is 1, like any other number.\[\frac{4}{7}=\theta\]

OpenStudy (mathmale):

Yes, bagajr has done a nice job of putting that into mathematical symbols. That's your desired answer. The angle, theta, is equal to 4/7 radians, which you could show is about 33 degrees.

OpenStudy (anonymous):